OpenStudy (itsmichelle29):

What is the exponential function modeled by the following table? x f(x) 2 10 3 28 4 82 f(x) = 2x f(x) = 2x + 1 f(x) = 3x f(x) = 3x + 1

OpenStudy (anonymous):

To this kind of problem, the fastest way is testing for example If f(x) = 2^x, Now, plug in x =2, f(x) =2^2 =4 , not =10 as given information --> f(x) =2^x is NOT the correct one. You do the rest, please

OpenStudy (itsmichelle29):

thanks